On January 1, 2021, HGC Camera Store adopted the dollar-value LIFO retail inventory method. Inventory transactions at both cost and retail, and cost indexes for 2021 and 2022 are as follows:

2021 2022
Cost Retail Cost Retail
Beginning inventory $ 57,000 $ 76,000
Net purchases 109,160 126,000 $ 115,150 $ 132,400
Freight-in 3,800 4,300
Net markups 19,000 11,600
Net markdowns 3,800 4,000
Net sales to customers 124,460 120,440
Sales to employees (net of 10% discount) 4,500 5,850
Price Index:
January 1, 2021 1.00
December 31, 2021 1.07
December 31, 2022 1.12


Required:
Estimate the 2021 and 2022 ending inventory and cost of goods sold using the dollar-value LIFO retail inventory method. (Do not round other intermediate calculations. Round your cost-to-retail percentage calculation to 2 decimal places and final answers to the nearest whole dollar.)
Answer #1


HGC Camera Store Estimate the 2021 and 2022 ending inventory and cost of goods sold using the dollar-value LIFO retail invent

For 2022: 5850 / 0.90 = 6500 - 5850 = 650 = Employee Discounts Beginning inventory Plus: Net Purchases Cost 62136 115150 4300